Hello, I need help with a problem. Use the commutative law to find a equivalent expression. An equivalant expression of uv is.
I think I should just reverse the order uv = vu is an equivalent expression.
Thanks for your time , I'm doing a quiz and my GPA in school is bad so I must ace all my assignments. I thank you if you have a little time to help.

Answer by Earlsdon(6294)

You are correct!
The "commutative law" says that, for addition and multiplication, the order in which the numbers or variables appear will make no difference to the result.
So...
5*8 = 8*5
23+7 = 7+12
This law does not apply to subtraction or division.